# High brightness multi-MeV photon source driven by a petawatt-scale laser wakefield accelerator

### v1

@@ Abstract
We present an experimental demonstration of a bright multi-MeV gamma source driven by a petawatt laser. The source generates on average $(1.2\pm 0.6)\times 10^{9}\,\mathrm{photons}$ above $1\,\mathrm{MeV}$ per pulse, exceeding those of previous all-optical sources by a hundred times, and reached a peak spectral brightness of $(3.9\pm 1.5)\times 10^{22}\,\mathrm{photons/mm^{2}/mrad^{2}/s/0.1\%BW}$ at $\epsilon_{\gamma}\approx 11\,\mathrm{MeV}$. The source was produced by inverse Compton scattering of a laser wakefield accelerated GeV electron beam and its back-reflected driving laser pulse, and is well described by a simple model of the laser and electron properties at the collision point. Our results highlight the promise of this source for fundamental physics studies, as well as for applications of nuclear resonance fluorescence and nuclear transmutation.
@@
All-optical ICS can also be produced with a single laser pulse, by using a plasma mirror to reflect the LWFA laser driver back onto the accelerated electron beam (see inset of Fig. 1), as demonstrated by Ta Phuoc et al. in 2012 [33]. This ‘self-aligned’ scheme guarantees alignment on every shot (a challenge for two-beam experiments [34, 35]) and essentially halves the laser requirements while enabling ICS sources at the more commonly existing single-beam laser facilities. Initial experiments have demonstrated this scheme at laser powers of 30 TW yielding up to 100 keV photon energies [33, 36, 37], extending up to MeV photon energies at the 100 TW scale [38, 39]. With this work, we extend the single-beam approach to the recently more widely available petawatt laser powers. Its high peak flux and ultrafast pulse duration provide novel capabilities that complement existing radiofrequency-based ICS sources, e.g. enabling imaging of fast processes or studies of fundamental physics such as photon-photon scattering [40]
@@
The distance from the tape to the exit of the gas jet nozzle was varied from 35 mm to 2.5 mm. To avoid damage to the laser system, the tape could not be placed closer than 2.5 mm as the gas affected the pointing stability too severely. The measured energy deposited in the gamma profiler $E_{dep}$ decreased with increasing distance between the tape and the gas jet (see Fig. 2), consistent with ICS as the laser pulse diffracts with distance, reducing its intensity at the interaction point. Other sources of radiation relevant in this context, e.g., bremsstrahlung or betatron radiation (resulting from oscillations of the electron beam in the plasma), do not depend on the position of the tape. The contribution of betatron radiation was characterised without the plasma mirror and was found to be negligible. Based on Geant4 simulations, the expected energy deposition in the gamma profiler from bremsstrahlung due to interaction of the electron beam with the tape ($\sim(2.3\pm 0.5)\times 10^{6}\,\mathrm{MeV}$) only becomes a sizeable fraction of the signal ($>10\%$) at the largest distance ($z=35\,\mathrm{mm}$), i.e. the signal is dominated by ICS for the measured range of distances.
Figure 2: Average energy deposited in the gamma profiler $E_{dep}$ (error bars for standard deviation) for increasing distance $z$ between the tape and the end of the gas jet. The emitted power and $E_{dep}$ are expected to scale with the intensity at the interaction point, which decreases with expanding beam area, i.e. $E_{dep}(z)=A[w_{0}/w(z)]^{2}+C$, where the Gaussian waist is $w(z)=w_{0}\sqrt{1+[(z-z_{0})/z_{R}]^{2}}$. The fitting parameters are the amplitude $A$, the constant offset $C$, the waist position $z_{0}$, and the Rayleigh length $z_{R}$. The fitting parameters for $z_{0}$ and $z_{R}$ are shown in the legend. The dashed line indicates $z_{0}$. Inset: Example gamma profiles at 2.5 mm (a), 5 mm (b) and 35 mm (c) distance, shown on the same scale. Counts on (c) were multiplied by 5 for visibility.
@@
While in the plasma, the laser drives a highly nonlinear plasma wave and experiences self-guiding, maintaining a high $a_{0}$, throughout the propagation [49]. After propagating to $z=25\,\mathrm{mm}$ (the closest collision point in the experiment), the laser had diffracted to a spot size of 70 µm fwhm and was trailed by an electron beam with $1.0\,\mathrm{nC}$ above 260 MeV up to a maximum energy of $1.5\,\mathrm{GeV}$ (Fig. 4b) and d)). The laser peak amplitude at this position was $a_{0}=2.0$ (4c) with an average wavelength of 1 µm due to redshifting in the plasma. The simulated electron beam had a transverse rms beam size of 4.3 µm, and a bunch length of 9.6 µm. For a wavelength of 1 µm, the Rayleigh length $z_{R}=\pi w_{0}^{2}/\lambda$ fitted to the experimental data in Fig. 2 corresponds to $w_{0}=27$ µm, consistent with the guided laser spot size in the simulation.
@@
By combining the electron beam distribution from the PIC simulations with the measured properties of the photon source, we can estimate the spectral brightness and compare its performance with other sources (see Fig. 5). The peak brightness, averaged over the 10 shots in Fig. 3, was $(0.9\pm 0.5)\times 10^{22}\,\mathrm{photons/mm^{2}/mrad^{2}/s/0.1\%BW}$ at a photon energy of 5 MeV. The single brightest shot had a peak brightness of $(3.9\pm 1.5)\times 10^{22}\,\mathrm{photons/mm^{2}/mrad^{2}/s/0.1\%BW}$ at 11 MeV. This is over 100 times higher than previous two-beam results at a similar photon energy [27], and has a similar peak brightness to previous single-beam results [38] but at 10 times higher photon energy. The peak brightness is just under half of the value reported from a two-beam experiment by Mirzaie et al. [30], which used about four times the laser power. However, the photon flux in our case is 100 times higher, with the lower peak brightness due to the comparably large divergence of the photon beam, inherited from that of the electron beam ($\approx 4$ mrad).
@@
where $W_{L}$ is the guided laser energy, $\tau/2$ is the interaction time in counterpropagating geometry and $\chi\approx 2\hbar\omega_{0}\gamma_{0}a_{0}/m_{e}c^{2}$. For $a_{g,0}=4$ (nominal value for the nonlinear bubble regime), this reduces to $W_{ICS}\approx 4\times 10^{-4}W_{L}^{2}$, where $\eta_{ICS}=4\times 10^{-4}W_{L}$ acts as an energy-dependent efficiency term. In the case of our experiment, the amount of laser energy participating in the LWFA was $W_{L}\approx 5.8$ J and so the expected ICS gamma beam energy is $W_{ICS}\approx 13$ mJ, in approximate agreement with the brightest experimental measurement. While ICS beam generation becomes more efficient at higher laser energy, further increasing the efficiency term could be achieved through increasing the laser intensity at the collision, for example, by increasing the plasma density close to the plasma mirror such that the laser self-focuses to a higher $a_{0}$.
@@
Figure 5: Spectral brightness of our work (orange) compared to other all-optical radiation sources using single-beam [33, 36, 38, 39] (green circle) or two-beam ICS [25, 26, 27, 28, 30] (blue triangle), and the SPring-8 synchrotron source (extracted from [55]). Lines indicate spectral brightness and markers peak spectral brightness. Radiofrequency-based ICS sources provide high average flux of order $10^{14}$ photons/s at lower peak brightness $<10^{15}$ [56, 9] up to MeV photon energies, at an average flux of $10^{6}-10^{10}$ photons/s at the MeV to GeV scale [7, 57, 58, 10]

### v2

@@ Abstract
We present an experimental demonstration of a bright multi-MeV gamma source driven by a petawatt laser. The source generates on average $(1.2\pm 0.6)\times 10^{9}\,\mathrm{photons}$ above $1\,\mathrm{MeV}$ per pulse, exceeding those of previous all-optical sources by a hundred times, and reached a peak spectral brightness of $(3.9\pm 1.5)\times 10^{22}\,\mathrm{photons/mm^{2}/mrad^{2}/s/0.1\%BW}$ at $\epsilon_{\gamma}\approx 11\,\mathrm{MeV}$. The source was produced by inverse Compton scattering of a laser wakefield accelerated GeV electron beam and its back-reflected driving laser pulse. Its performance is well described by a simple model of the laser and electron properties at the collision point that allows quantitative predictions and identifies clear strategies to further enhance radiation efficiency. Our results highlight the promise of this source for fundamental physics studies, as well as for applications of nuclear resonance fluorescence and nuclear transmutation.
@@
All-optical ICS can also be produced with a single laser pulse, by using a plasma mirror to reflect the LWFA laser driver back onto the accelerated electron beam (see inset of Fig. 1), as demonstrated by Ta Phuoc et al. in 2012 [33]. This ‘self-aligned’ scheme guarantees alignment on every shot (a challenge for two-beam experiments [34, 35]) and essentially halves the laser requirements while enabling ICS sources at the more commonly existing single-beam laser facilities. For practical, high-impact applications, it is critical to optimise the ICS efficiency to maximise photon yield with limited laser energy. Initial experiments have demonstrated this scheme at laser powers of 30 TW yielding up to 100 keV photon energies [33, 36, 37], extending up to MeV photon energies at the 100 TW scale [38, 39]. With this work, we extend the single-beam approach to the recently more widely available petawatt laser powers. By generating photons with energies exceeding 10 MeV, the source becomes directly relevant to nuclear physics applications and substantially broadens the range of potential uses. Its high peak flux and ultrafast pulse duration provide novel capabilities that complement existing radiofrequency-based ICS sources, e.g. enabling imaging of fast processes or studies of fundamental physics such as photon-photon scattering [40].
@@
The distance from the tape to the exit of the gas jet nozzle was varied from 35 mm to 2.5 mm. To avoid damage to the laser system, the tape could not be placed closer than 2.5 mm as the gas affected the pointing stability too severely. The measured energy deposited in the gamma profiler $E_{dep}$ decreased with increasing distance between the tape and the gas jet (see Fig. 2), consistent with ICS as the laser pulse diffracts with distance, reducing its intensity at the interaction point. Other sources of radiation relevant in this context, e.g., bremsstrahlung or betatron radiation (resulting from oscillations of the electron beam in the plasma), do not depend on the position of the tape. The background signal, e.g. caused by betatron radiation or bremsstrahlung from electrons striking the vacuum chamber walls, was characterised without the plasma mirror and was found to be comparable to the signal for $z>15\,\mathrm{mm}$ ($\sim(4.1\pm 0.1)\times 10^{7}$ MeV). Based on Geant4 simulations, the expected energy deposition in the gamma profiler from bremsstrahlung due to interaction of the electron beam with the tape was found to be negligible ($\sim(3.3\pm 0.6)\times 10^{6}\,\mathrm{MeV}$). The signal was dominated by ICS for the closest distances ($z<5\,\mathrm{mm}$).
Figure 2: Average energy deposited in the gamma profiler $E_{dep}$ (error bars for standard deviation) for increasing distance $z$ between the tape and the end of the gas jet. The emitted power and $E_{dep}$ are expected to scale with the intensity at the interaction point, which decreases with expanding beam area, i.e. $E_{dep}(z)=A[w_{0}/w(z)]^{2}+C$, where the Gaussian waist is $w(z)=w_{0}\sqrt{1+[(z-z_{0})/z_{R}]^{2}}$. The fitting parameters are the amplitude AA, the constant offset $C$, the waist position $z_{0}$, and the Rayleigh length $z_{R}$. The fitting parameters for $z_{0}$ and $z_{R}$ are shown in the legend. The dashed line indicates $z_{0}$. Inset: Example gamma profiles at 2.5 mm (a), 5 mm (b) and 35 mm (c) distance, shown on the same scale. Counts on (c) were multiplied by 5 for visibility.
@@
While in the plasma, the laser drives a highly nonlinear plasma wave and experiences self-guiding, maintaining a high $a_{0}$, throughout the propagation [49]. After propagating to $z=25\,\mathrm{mm}$ (the closest collision point in the experiment), the laser had diffracted to a spot size of 70 µm fwhm and was trailed by an electron beam with $1.0\,\mathrm{nC}$ above 260 MeV up to a maximum energy of $1.5\,\mathrm{GeV}$ (Fig. 4b) and d)). The laser peak amplitude at this position was $a_{0}=2.0$ (4c) with an average wavelength of 1 µm due to redshifting in the plasma. For a wavelength of 1 µm, the Rayleigh length $z_{R}=\pi w_{0}^{2}/\lambda$ fitted to the experimental data in Fig. 2 corresponds to $w_{0}=27$ µm, consistent with the guided laser spot size in the simulation. The simulated electron beam had a transverse rms beam size of 4.3 µm, and a bunch length of 9.6 µm.
@@
By combining the electron beam distribution from the PIC simulations with the measured properties of the photon source, we can estimate the spectral brightness and compare its performance with other sources (see Fig. 5). The peak brightness, averaged over the 10 shots in Fig. 3, was $(0.9\pm 0.5)\times 10^{22}\,\mathrm{photons/mm^{2}/mrad^{2}/s/0.1\%BW}$ at a photon energy of 5 MeV. The single brightest shot had a peak brightness of $(3.9\pm 1.5)\times 10^{22}\,\mathrm{photons/mm^{2}/mrad^{2}/s/0.1\%BW}$ at 11 MeV. This is over 100 times higher than previous two-beam results at a similar photon energy [27], and has a similar peak brightness to previous single-beam results [38] but at 10 times higher photon energy. The peak brightness is just under half of the value reported from a two-beam experiment by Mirzaie et al. [30], which used about four times the laser power. However, the photon flux in our case is 100 times higher and collisions are observed on every shot due to the self-aligning geometry. The high photon yield is enabled by the interaction with a high-charge electron beam ($\approx 1\,\mathrm{nC}$) and improved scattering efficiency that maximises the participating charge. In contrast, in the experiment of Mirzaie et al. [30], less than 1% of electrons were overlapped with the laser pulse to optimise collision probability. Importantly, for many practical applications, high photon flux, reduced laser requirements, and consistent operation are critical performance metrics, making this approach particularly attractive. The lower peak brightness observed here is primarily due to the comparably large divergence of the photon beam, which is inherited from the divergence of the electron beam ($\approx 4$ mrad).
@@
where $W_{L}$ is the guided laser energy, $\tau/2$ is the interaction time in counterpropagating geometry and $\chi\approx 2\hbar\omega_{0}\gamma a_{0}/m_{e}c^{2}$. For $a_{g,0}=4$ (nominal value for the nonlinear bubble regime), this reduces to $W_{ICS}\approx{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\eta_{ICS}W_{L}\mathrm{[J]}}}$, where $\eta_{ICS}=4\times 10^{-4}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\mathrm{[J^{-1}}]}}W_{L}$ acts as a dimensionless energy-dependent efficiency term. In the case of our experiment, the amount of laser energy participating in the LWFA was $W_{L}\approx 5.8$ J and so the expected ICS gamma beam energy is $W_{ICS}\approx 13$ mJ, in approximate agreement with the brightest experimental measurement of $21\pm 9\,\mathrm{mJ}$. While ICS beam generation becomes more efficient at higher laser energy, further increasing the efficiency term could be achieved through increasing the laser intensity at the collision, for example, by increasing the plasma density close to the plasma mirror such that the laser self-focuses to a higher $a_{0}$.
@@
Figure 5: Spectral brightness of our work (orange) compared to other all-optical radiation sources using single-beam [33, 36, 38, 39] (green circle) or two-beam ICS [25, 26, 27, 28, 30] (blue triangle), and the SPring-8 synchrotron source (extracted from [55]). Lines indicate spectral brightness and markers peak spectral brightness. Radiofrequency-based ICS sources provide high average flux of order $10^{14}$ photons/s at lower peak brightness $10^{15}$ [56, 9] up to MeV photon energies, at an average flux of $10^{6}-10^{10}$ photons/s at the MeV to GeV scale [7, 57, 58, 10]